# Learned Free-Energy Functionals from Pair-Correlation Matching for Dynamical Density Functional Theory

### v2

@@ V.1 Brownian dynamics
We use LAMMPS [29] to simulate BD motion of Lennard-Jones (LJ) particles under repulsive external fields. The motion is specified by the following overdamped Langevin equation,
$d\mathbf{r}=\gamma^{-1}\mathbf{F}\,dt+\sqrt{2k_{B}T\,\gamma^{-1}}\,d\mathbf{W}_{t},$ (12)
@@
We run BD simulations inside a closed $3$D cube box of edge length $10\sigma$ with periodic boundary conditions on all axes. We randomly initialize $600$ LJ particles (with a cut-off distance of $4\epsilon$) inside the box. $\epsilon$ and $\sigma$ are the LJ coefficients of well-depth and particle diameter respectively. The temperature of the system is kept at a constant $k_{B}T/\epsilon=2$ and each timestep is $10^{-3}\tau$ where $\tau$ is the time scale $\gamma\sigma^{2}/\epsilon$ and $\gamma$ is the mobility coefficient.
@@ V.2 Dynamical density functional theory
We numerically integrate the DDFT equation (section III), with the neural excess free-energy functional from pair-correlation matching,
@@
to obtain time-dependent density profiles of $3$D Lennard-Jones (LJ) systems with planar geometry under an external potential $V_{\text{ext}}$ (Fig. 1).
We define a $1$D density grid of $10\sigma$ length with a spacing of $\sigma/100$ for the solver, with periodic boundaries. We initialize the density grid with the same uniform initial density as the reference BD simulation and apply the same external potential. We simulate the PDE with a time-step of $10^{-4}\tau$ until $t=4\tau$, where $\tau=\sqrt{{m\sigma^{2}}/{\epsilon}}$ in reduced LJ units. The gradients are computed using finite-differences and time integrated using an Euler scheme, as implemented in the numerical Py-PDE Python package [31]. A psuedospectral solver can also be used for scenarios that require more numerical precision [32]. The mobility coefficient $\Gamma$ is chosen to be the same as in BD. For computing ${\delta\mathcal{F}_{\theta,\text{exc}}[\rho(\mathbf{r},t)]}/{\delta\rho(\mathbf{r},t)}$, we use automatic differentiation over the excess neural free-energy functional.
@@ VI Experiments
Figure 4: Breakthrough simulations using neural DDFT and Brownian dynamics (see Appendix E for the movie version). (a) LJ particles from the source move towards the sink under two different MOF-like repulsive external potentials. The dynamics of the density evolution and the final evolved density (visualized) obtained from DDFT are in close agreement with BD. (b) Time evolution of the output to input flux ratio, showing the approach towards steady-state or breakthrough condition. The particles reach breakthrough earlier under the MOF-2 external potential. The curves obtained from DDFT are in close agreement with BD, while being smoother and taking only $2\%$ of the compute time of BD.
@@ VI.1 Closed system
Our experiments show that the neural free energy functionals $\mathcal{F}_{\theta,\text{exc}}^{(1)}$ and $\mathcal{F}_{\theta,\text{exc}}^{(2)}$ are both significantly better than the $\mathcal{F}^{\text{MF}}_{\text{exc}}$ functional in accuracy across time, as shown in. Fig. 2 & Fig. 3. The convergence of $\mathcal{F}_{\theta,\text{exc}}^{(2)}$ is similar to $\mathcal{F}_{\theta,\text{exc}}^{(1)}$, except in Fig. 3(d) where the density near the wells is pushed above the maximum bulk value of $0.67\sigma^{-3}$ seen during training of $\mathcal{F}_{\theta,\text{exc}}^{(2)}$. The errors for all the DDFT approaches for time intervals $t<1\tau$ are high which we attribute to overly fast dynamics arising from the adiabatic approximation [22]. More notably, the dynamical density profiles from DDFT are several orders of magnitude faster to compute with DDFT than with BD simulation: while running $5000$ trials of BD takes about a day on a single-core CPU, the DDFT approach on the same computer takes less than 6 minutes (see Appendix D for runtime comparisions).
@@ VI.2 Open system
To evaluate the pair-correlation matching functional $\mathcal{F}_{\theta,\text{exc}}^{(2)}$ on an open system, we construct a system modeled after “breakthrough” experiments used in gas adsorption studies [34, 35]. In these studies, typically a column filled with adsorbent material (e.g. nanoporous metal-organic framework (MOF) is exposed to a gas stream (e.g. CO2), where the breakthrough of the gas is detected at the outlet to estimate the adsorption capacity and kinetics of the material. We adopt such a setup by extending the closed Lennard-Jones system from the previous experiment with two particle baths, one at a high $\mu=0.5$ in $z\in[0,2]\sigma$ to simulate an LJ particle source and the other at extremely low $\mu=-100$ in $z\in[8,10]\sigma$ to simulate an LJ particle sink. The difference in $\mu$ causes particles to flow from the source to the sink through the region $z\in[2,8]\sigma$, in which we apply, in addition, a planar potential to mimic two MOF planes from the IRMOF series [36] (see Appendix C). In the reference BD simulation, we extend the closed system to an open system by attaching two Grand-Canonical Monte-Carlo (GCMC) reservoirs as source and sink. The source reservoir is from $z=[0,2]\sigma$ with $\mu=0.5$ and the sink is from $z=[8,10]\sigma$ with $\mu=-100$. 10 trials of GCMC insertion/deletion steps, with no translation/rotation, are attempted in LAMMPS (hybrid) between every BD update. Before starting the breakthrough simulation, we first randomly initialize LJ particles in the source region and confine them using a potential barrier and equilibrate the system for $20000$ steps (each timestep is $10^{-3}\tau$). Then the barrier is removed and the MOF-like potential is applied on the system and $100000$ collection steps are performed as the particles flow towards the sink. The planar density is sampled every 100 steps, over a bin size of ${\sigma}/{32}$ and is averaged over $500$ BD runs.
We simulate this same system using our open DDFT equation (eq. 8) for a duration of $150\tau$, with a timestep of $10^{-3}\tau$ and grid spacing of ${\sigma}/{32}$. To obtain breakthrough curves we compute the time-dependence of the ratio $J(z_{2},t)/J(z_{1},T)$ of the particle fluxes $J(z_{2},t)$ close to the output at $z_{2}=7.5\sigma$ and $J(z_{1},T)$ close to the input at $z_{1}=2.5\sigma$ at late time $T=145\tau$ where steady-state is (almost) reached. The particle flux from DDFT is computed as $J(z,t)=-\Gamma\rho(z,t)\nabla\left({\delta\mathcal{F}[\rho(z)]}/{\delta\rho(z)}\right)$, where $\mathcal{F}_{\theta,\text{exc}}^{(2)}$ is used for the excess component. We also apply a Savitsky-Golay filter [37] over the spatial gradient in the flux computation to smooth out any noise arising from numerical errors. The BD flux is computed as the net number of particles crossing the planes at $z_{1}$ and $z_{2}$ per unit time per unit area, averaged over $500$ BD runs.
@@ VII Conclusion
To summarize, we have presented results that the neural excess free-energy functional trained via pair-correlation matching can be directly applied to DDFT without retraining, with high accuracy relative to an analytical free-energy approximation and with high efficiency relative to BD simulations. We show this for both conserved dynamics, as well as for non-conserved dynamics via a novel gradient flow based extension which we evaluate on a system inspired by breakthrough adsorption simulations. Future work could focus on the extension of this method to systems with full $3$D geometry and anisotropic particles, and extending it to incorporate superadiabatic forces [38], towards practical applications in modeling many-body non-equilibrium systems.
@@ Acknowledgements
We would like to thank David Dubbeldam and Christian Koke for useful discussions.
@@
Code and data to reproduce all results will be made available upon publication.
@@ Appendix A Classical Density Functional Theory
The challenge is to approximate $\mathcal{F}_{\text{exc}}[\rho]$ which represents the inter-particle interactions as a functional of the non-local density. Once such an $\mathcal{F}[\rho]$ is (approximately) obtained, the equilibrium density $\rho_{\text{eq}}(\rho)$ from Eq. 14 can be formulated as the following Euler-Lagrange equation,
@@ A.1 Single-Body Direct Correlation Matching
In addition to pair-correlation matching [10], we can also train a neural free-energy functional ($\mathcal{F}_{\theta,\text{exc}}^{(1)}$) using single-body direct correlation matching. This is done by minimizing the error between the target ${\delta\mathcal{F}_{\text{exc}}}/{\delta\rho(z_{i})}$ and the estimate $(1/{\Delta z}){\partial\mathcal{F}_{\theta,\text{exc}}^{(1)}}/{\partial\rho_{i}}$. This functional derivative of the excess free-energy is the one-body direct correlation function, and computing the target samples requires simulating inhomogeneous equilibrium densities under a range of external potentials.
@@ Appendix B Dynamical Density Functional Theory
$\frac{\partial\rho(\mathbf{r},t)}{\partial t}+\nabla\cdot\mathbf{J}(\mathbf{r},t)=0.\$ (18)
@@
$\frac{\partial\rho(\mathbf{r},t)}{\partial t}+\nabla\cdot\left(\rho(\mathbf{r},t)\mathbf{v}(\mathbf{r},t)\right)=0,\$ (19)
@@
$\frac{\partial\rho(\mathbf{r},t)}{\partial t}-\Gamma\nabla\cdot\left(\rho(\mathbf{r},t)\nabla\mu(\mathbf{r},t)\right)=0.\$ (20)
@@
Re-arranging the terms and expanding the free-energy functional we get,
@@ Appendix C MOF-like planar potentials
Here we describe our procedure for obtaining the $1$D external potentials from metal-organic frameworks (MOF) for the breakthrough simulation experiment in the main paper.
We use the IRMOF-1 and IRMOF-10 structures from the IRMOF series of crystal structures [36], which have the same topology but differing pore sizes. To convert their 3D structures into effective 1D external potentials suitable for DDFT, we employed a simulation-based force sampling approach using LAMMPS [29]. Starting from CIF files of the MOFs, we generated compatible data files for LAMMPS with appropriate force field assignments and topology using lammps-interface [40]. The pore axis of each MOF was aligned along the z-direction, and a single LJ-type guest particle was introduced and displaced incrementally along its pore axis and the total force acting on it was recorded at each step. This produced a raw force profile as a function of z-position which was then converted to reduced LJ units and numerically integrated (in reverse) to generate a dimensionless potential profile. The resulting potentials were clipped and centered to remove artifacts at the boundaries and padded with zeros to enforce a flat, finite domain. The final 1D potentials, expressed as functions interpolated over a grid, serve as external fields $V_{\text{ext}}$ that approximate the confinement environment of the original MOF pores.
@@ Appendix D Runtime comparisons
The below table is a comparison of approximate runtimes between BD simulation and our DDFT implementation for open and closed systems across two different grid sizes. While each DDFT simulation directly produces smooth density profiles, each BD simulation needs to be ensemble averaged (over 5000 and 500 runs for closed and open resp.), to obtain comparable density profiles. Both simulations were run on a single-core of Intel(R) Xeon(R) W-2123 CPU @ 3.60GHz or comparable CPU. Note that DDFT uses $\Delta t=10^{-3}\tau$ for grid spacing ${\sigma}/{32}$ and $\Delta t=10^{-4}\tau$ for grid spacing ${\sigma}/{100}$, while BD uses $\Delta t=10^{-3}\tau$ for both.
@@
Closed (Eq. (M1), ${\sigma}/{32}$ 34 hours 25 secs.
until $t=4\tau$) ${\sigma}/{100}$ 34 hours 5 mins.
Open (Eq. (M2), ${\sigma}/{32}$ 48 hours 40 mins.
until $t=150\tau$) ${\sigma}/{100}$ 48 hours 8 hours
Appendix E Supplemental Movies
We also attach two movie files as supplemental material (also available on the project webpage [41]).
dynamical-density-profiles.mp4: Dynamical version of Fig. (3) showing the density profiles in the main paper.
breakthrough-simulation.mp4: Dynamical version of Fig. (4) showing the breakthrough curves in the main paper.

### v3

@@ V.1 Brownian dynamics
We use LAMMPS [30] to simulate BD motion of Lennard-Jones (LJ) particles under repulsive external fields. The motion is specified by the following overdamped Langevin equation:
$d\mathbf{r}=\Gamma\mathbf{F}\,dt+\sqrt{2k_{B}T\,\Gamma}\,d\mathbf{W}_{t},$ (12)
@@
We run BD simulations inside a closed 33D cube box of edge length $10\sigma$ with periodic boundary conditions on all axes. We randomly initialize $600$ LJ particles (with a cut-off distance of $4\epsilon$) inside the box. $\epsilon$ and $\sigma$ are the LJ coefficients of well-depth and particle diameter respectively. The temperature of the system is kept at a constant $k_{B}T/\epsilon=2$ and each timestep is $10^{-3}\tau$ where $\tau$ is the time scale $(\sqrt{{m\sigma^{2}}/{\epsilon}})$.
@@ V.2 Dynamical density functional theory
We numerically integrate the DDFT equation [section III], with the neural excess free-energy functional from pair-correlation matching,
@@
to obtain time-dependent density profiles of 33D Lennard-Jones (LJ) systems with planar geometry under an external potential $V_{\text{ext}}(\mathbf{r})$ (Fig. 1).
We define a 11D density grid of $10\sigma$ length with a spacing of $\sigma/100$ for the solver, with periodic boundaries. We initialize the density grid with the same uniform initial density as the reference BD simulation and apply the same external potential. We simulate the PDE with a time-step of $10^{-4}\tau$ until $t=4\tau$, where $\tau=\sqrt{{m\sigma^{2}}/{\epsilon}}$ in reduced LJ units. The gradients are computed using finite-differences and time integrated using an Euler scheme, as implemented in the numerical Py-PDE Python package [32]. A pseudospectral solver can also be used for scenarios that require more numerical precision [33]. The mobility coefficient $\Gamma$ is chosen to be the same as in BD. For computing ${\delta\mathcal{F}_{\theta,\text{exc}}[\rho(\mathbf{r},t)]}/{\delta\rho(\mathbf{r},t)}$, we use automatic differentiation over the excess neural free-energy functional.
@@ VI Experiments
Figure 4: Breakthrough simulations using neural DDFT and Brownian dynamics (see SM [29] for the movie version). (a) LJ particles from the source move towards the sink under two different MOF-like repulsive external potentials. The dynamics of the density evolution and the final evolved density (visualized) obtained from DDFT are in close agreement with BD. (b) Time evolution of the output to input flux ratio, showing the approach towards steady-state or breakthrough condition. The particles reach breakthrough earlier under the MOF-2 external potential. The curves obtained from DDFT are in close agreement with BD, while being smoother and taking only $2\%$ of the compute time of BD.
@@ VI.1 Closed system
Our experiments show that the neural free energy functionals $\mathcal{F}_{\theta,\text{exc}}^{(1)}$ and $\mathcal{F}_{\theta,\text{exc}}^{(2)}$ are both significantly better than the $\mathcal{F}^{\text{MF}}_{\text{exc}}$ functional in accuracy across time, as shown in Fig. 2 and Fig. 3. The convergence of $\mathcal{F}_{\theta,\text{exc}}^{(2)}$ is similar to $\mathcal{F}_{\theta,\text{exc}}^{(1)}$, except in Fig. 3(d) where the density near the wells is pushed above the maximum bulk value of $0.67\sigma^{-3}$ seen during training of $\mathcal{F}_{\theta,\text{exc}}^{(2)}$. The errors for all the DDFT approaches for time intervals $t<1\tau$ are high which we attribute to overly fast dynamics arising from the adiabatic approximation [22]. More notably, the dynamical density profiles from DDFT are several orders of magnitude faster to compute with DDFT than with BD simulation: while running $5000$ trials of BD takes about a day on a single-core CPU, the DDFT approach on the same computer takes less than 6 minutes (see Appendix D for runtime comparisons).
@@ VI.2 Open system
To evaluate the pair-correlation matching functional $\mathcal{F}_{\theta,\text{exc}}^{(2)}$ on an open system, we construct a system modeled after “breakthrough” experiments used in gas adsorption studies [35, 36]. In these studies, typically a column filled with adsorbent material [e.g. nanoporous metal-organic framework (MOF)] is exposed to a gas stream (e.g. CO2), where the breakthrough of the gas is detected at the outlet to estimate the adsorption capacity and kinetics of the material. We adopt such a setup by extending the closed Lennard-Jones system from the previous experiment with two particle baths, one at a high $\beta\mu=0.5$ in $z\in[0,2]\sigma$ to simulate an LJ particle source and the other at extremely low $\beta\mu=-100$ in $z\in[8,10]\sigma$ to simulate an LJ particle sink. The difference in $\mu$ causes particles to flow from the source to the sink through the region $z\in[2,8]\sigma$, in which we apply, in addition, a planar potential to mimic two MOF planes from the IRMOF series [37] (see Appendix C). In the reference BD simulation, we extend the closed system to an open system by attaching two Grand-Canonical Monte-Carlo (GCMC) reservoirs as source and sink. The source reservoir is from $z=[0,2]\sigma$ with $\beta\mu=0.5$ and the sink is from $z=[8,10]\sigma$ with $\beta\mu=-100$. Ten trials of GCMC insertion/deletion steps, with no translation/rotation, are attempted in LAMMPS (hybrid) between every BD update. Before starting the breakthrough simulation, we first randomly initialize LJ particles in the source region and confine them using a potential barrier and equilibrate the system for $20000$ steps (each timestep is $10^{-3}\tau$). Then the barrier is removed and the MOF-like potential is applied on the system and $100000$ collection steps are performed as the particles flow towards the sink. The planar density is sampled every 100 steps, over a bin size of ${\sigma}/{32}$ and is averaged over $500$ BD runs.
We simulate this same system using our open DDFT equation [eq. 8] for a duration of $150\tau$, with a timestep of $10^{-3}\tau$, grid spacing of ${\sigma}/{32}$, and $\Gamma^{\prime}=1$. To obtain breakthrough curves we compute the time-dependence of the ratio $J(z_{2},t)/J(z_{1},T)$ of the particle fluxes $J(z_{2},t)$ close to the output at $z_{2}=7.5\sigma$ and $J(z_{1},T)$ close to the input at $z_{1}=2.5\sigma$ at late time $T=145\tau$ where steady-state is (almost) reached. The particle flux from DDFT is computed as $J(z,t)=-\Gamma\rho(z,t)\nabla\left({\delta\mathcal{F}[\rho(z)]}/{\delta\rho(z)}\right)$, where $\mathcal{F}_{\theta,\text{exc}}^{(2)}$ is used for the excess component. We also apply a Savitsky-Golay filter [38] over the spatial gradient in the flux computation to smooth out any noise arising from numerical errors. The BD flux is computed as the net number of particles crossing the planes at $z_{1}$ and $z_{2}$ per unit time per unit area, averaged over $500$ BD runs.
@@ VII Conclusion
To summarize, we have presented results that the neural excess free-energy functional trained via pair-correlation matching can be directly applied to DDFT without retraining, with high accuracy relative to an analytical free-energy approximation and with high efficiency relative to BD simulations. We show this for both conserved dynamics, as well as for non-conserved dynamics via a novel gradient flow based extension which we evaluate on a system inspired by breakthrough adsorption simulations. Pair-correlation matching thus offers a practical route to predicting non-equilibrium dynamics by training on homogeneous equilibrium density data, which is significantly less expensive to sample in higher dimensions than inhomogeneous equilibrium density data [10], inhomogeneous steady-state density data [39] or out-of-equilibrium dynamical density data.
Future work could focus on extending this method to incorporate superadiabatic forces [40]. In this work, we relied on the adiabatic approximation of DDFT, i.e. the non-equilibrium internal force field can be approximated as a conservative force given by the gradient of the equilibrium single body direct correlation function, to drive the non-equilibrium dynamics. Non-conservative forces can be important, especially for shear flows, which requires learning a correction term to this equilibrium functional. However the main challenge in training such a correction term is that it would require training on dynamical trajectories from BD, which can be very expensive to simulate relative to equilibrium densities from GCMC. To address this, Zimmermann et al[39] recently proposed training a neural force functional on steady-state dynamical data, which are relatively easier to sample than non-equilibrium dynamical data, showing promising results in extrapolating to out-of-steady-state non-equilibrium dynamics with non-conservative forces. Other future work includes extensions to systems with full 33D geometry, anisotropic particles, and mixtures towards practical applications in modeling many-body non-equilibrium systems.
@@ Acknowledgements
We would like to thank David Dubbeldam and Christian Koke for useful discussions. J.D. and B.E. would like to thank the University of Amsterdam Data Science Centre for financial support. K.R. acknowledges his affiliation with the ELLIS (European Laboratory for Learning and Intelligent Systems) PhD program.
@@
The source code for BD and DDFT simulations, along with the model weights and resulting trajectories are made available at [41].
@@ Appendix A Classical Density Functional Theory
The challenge is to approximate $\mathcal{F}_{\text{exc}}[\rho]$ which represents the interparticle interactions as a functional of the non-local density. Once such an $\mathcal{F}[\rho]$ is (approximately) obtained, the equilibrium density $\rho_{\text{eq}}(\rho)$ from Eq. 14 can be formulated as the following Euler-Lagrange equation,
@@ A.1 Single-body direct correlation matching
In addition to pair-correlation matching [10], we can also train a neural free-energy functional ($\mathcal{F}_{\theta,\text{exc}}^{(1)}$) using single-body direct correlation matching (Fig. 5). This is done by minimizing the error between the target ${\delta\mathcal{F}_{\text{exc}}}/{\delta\rho(z_{i})}$ and the estimate $(1/{\Delta z}){\partial\mathcal{F}_{\theta,\text{exc}}^{(1)}}/{\partial\rho_{i}}$. This functional derivative of the excess free-energy is the one-body direct correlation function, and computing the target samples requires simulating inhomogeneous equilibrium densities under a range of external potentials.
@@ Appendix B Dynamical Density Functional Theory
$\frac{\partial\rho(\mathbf{r},t)}{\partial t}+\nabla\cdot\mathbf{J}(\mathbf{r},t)=0.\\$ (18)
@@
$\frac{\partial\rho(\mathbf{r},t)}{\partial t}+\nabla\cdot\left(\rho(\mathbf{r},t)\mathbf{v}(\mathbf{r},t)\right)=0,\\$ (19)
@@
$\frac{\partial\rho(\mathbf{r},t)}{\partial t}-\Gamma\nabla\cdot\left(\rho(\mathbf{r},t)\nabla\mu(\mathbf{r},t)\right)=0.\\$ (20)
@@
Re-arranging the terms and expanding the free-energy functional we obtain,
@@ Appendix C MOF-like planar potentials
Here we describe our procedure for obtaining the 11D external potentials from metal-organic frameworks (MOF) for the breakthrough simulation experiment in the main paper.
We use the IRMOF-1 and IRMOF-10 structures (Fig. 6) from the IRMOF series of crystal structures [37], which have the same topology but differing pore sizes. To convert their 3D structures into effective 1D external potentials suitable for DDFT, we employed a simulation-based force sampling approach using LAMMPS [30]. Starting from CIF files of the MOFs, we generated compatible data files for LAMMPS with appropriate force field assignments and topology using lammps-interface [43]. The pore axis of each MOF was aligned along the z-direction, and a single LJ-type guest particle was introduced and displaced incrementally along its pore axis and the total force acting on it was recorded at each step. This produced a raw force profile as a function of z-position which was then converted to reduced LJ units and numerically integrated (in reverse) to generate a dimensionless potential profile. The resulting potentials were clipped and centered to remove artifacts at the boundaries and padded with zeros to enforce a flat, finite domain. The final 1D potentials, expressed as functions interpolated over a grid, serve as external fields $V_{\text{ext}}(\mathbf{r})$ that approximate the confinement environment of the original MOF pores.
@@ Appendix D Runtime comparisons
The table below is a comparison of approximate runtimes between BD simulation and our DDFT implementation for open and closed systems across two different grid sizes. While each DDFT simulation directly produces smooth density profiles, each BD simulation needs to be ensemble averaged (over 5000 and 500 runs for closed and open, respectively), to obtain comparable density profiles. Both simulations were run on a single-core of Intel(R) Xeon(R) W-2123 CPU @ 3.60GHz or comparable CPU. Note that DDFT uses $\Delta t=10^{-3}\tau$ for grid spacing ${\sigma}/{32}$ and $\Delta t=10^{-4}\tau$ for grid spacing ${\sigma}/{100}$, while BD uses $\Delta t=10^{-3}\tau$ for both.
@@
Closed [Eq. (M1), ${\sigma}/{32}$ 34 h 25 s
until $t=4\tau$] ${\sigma}/{100}$ 34 h 5 min
Open [Eq. (M2), ${\sigma}/{32}$ 48 h 40 min
until $t=150\tau$] ${\sigma}/{100}$ 48 h 8 h